OpenStudy (anonymous):

An expression is shown below: 3x3y + 12xy - 9x2y - 36y Part A: Rewrite the expression so that the GCF is factored completely. Show the steps of your work. (3 points) Part B: Rewrite the expression completely factored. Show the steps of your work. (4 points) Part C: If the two middle terms were switched so that the expression became 3x3y - 9x2y + 12xy - 36y, would the factored expression no longer be equivalent to your answer in part B? Explain your reasoning. (3 points)

OpenStudy (anonymous):

help please

OpenStudy (anonymous):

desmond siaw

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

is this sanchez

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

do your own work bih

OpenStudy (anonymous):

lol help me

OpenStudy (anonymous):

give me a met

OpenStudy (anonymous):

metal

OpenStudy (anonymous):

ill give you one too

OpenStudy (anonymous):

lol

OpenStudy (amoodarya):

\[3x^3y + 12xy - 9x^2y - 36y=\\(3x^3y + 12xy )+(- 9x^2y - 36y)=\\3xy(x^2+4)-9y(x^2+4)=\\(x^2+4) (3xy-9y) =\\3y(x^2+4)(3-3)\]

OpenStudy (anonymous):

so thats the anser mr.

OpenStudy (anonymous):

part a